# REINFORCE++: An Efficient RLHF Algorithm with Robustness to Both Prompt and Reward Models

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) plays a crucial role in aligning large language models (LLMs) with human values and preferences. While state-of-the-art applications like ChatGPT/GPT-4 commonly employ Proximal Policy Optimization (PPO), the inclusion of a critic network introduces significant computational overhead. REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network. However, these approaches face challenges in accurate advantage estimation. Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on simpler prompts and vulnerability to reward hacking. To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the normalized reward of a batch as the baseline. Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation. Furthermore, it achieves superior generalization in both RLHF and long chain-of-thought (CoT) settings compared to existing REINFORCE-based methods. The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
@@ 1 Introduction
Without the critic network, REINFORCE-based methods often struggle to estimate individual tokens’ advantages accurately. Various REINFORCE-baseline approaches have been proposed to address this limitation, but each has significant drawbacks. ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation. RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO utilizes a normalized reward across all responses. These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking. Furthermore, these methods calculate reward baselines separately for each prompt, leading to overfitting and instability on specific training prompts during optimization. Therefore, these methods require carefully curating prompt sets particular to each task.
@@ 3 Method
REINFORCE++ still optimizes the PPO objective and employs the clipping strategy as defined in Equation 1. To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward. Accordingly, the advantage of REINFORCE++ is the normalized reward-to-go in reinforcement learning:
@@ 4.2.2 RL from Supervised Fine-tuned Model
In real-world tasks, user prompts exhibit enormous form and intent, making it difficult to assess a given model’s strengths and weaknesses systematically. Using synthetic datasets, we can systematically manipulate key factors such as length and difficulty, allowing for a more direct and interpretable evaluation and analysis of model performance.
@@ 5 Conclusion
In this paper, we presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences. Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global batch mean reward as a baseline to prevent overfitting and instability. Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities. Future work includes exploring adaptive normalization techniques, advanced variance reduction methods, and extending REINFORCE++ beyond RLHF settings.
@@ Appendix A Algorithm Details
Our experiments above showed that training models using multiple generated responses per prompt can further improve performance. Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation. Specifically, we sample multiple responses for each prompt and compute their average reward as the baseline to reshape the rewards. The reward for each prompt-response pair is then defined as:
$Adv_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(o_{t})$
$Adv_{q,o_{t}}^{norm}=\frac{Adv_{q,o_{t}}}{\text{std}_{batch}(Adv_{q,o_{t}})}$
@@
REINFORCE++ shares certain similarities with PPO in its formulation. Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and additionally employs global batch normalization as the baseline. Mathematically, this relationship is expressed as follows:
@@
where $V(s_{t})$ denotes the estimated value function at state $s_{t}$. By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term, while the introduction of global batch normalization further stabilizes training by reducing variance and preventing overfitting to individual prompts.
@@ Appendix B Acknowledgements
  • Jian Hu: Developed the ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the early draft of the paper.
@@
  • Shen Wei: Led the paper writing and designed and supervised the main experiments.

### v3

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) is crucial in aligning large language models (LLMs) with human values and preferences. While state-of-the-art applications like ChatGPT/GPT-4 commonly employ Proximal Policy Optimization (PPO), including a critic network introduces significant computational overhead. REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network. However, these approaches face challenges in accurate advantage estimation. Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on more straightforward prompts and vulnerability to reward hacking. To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the normalized reward of a batch as the baseline. Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation. Furthermore, it achieves superior generalization in RLHF and long chain-of-thought (CoT) settings compared to REINFORCE-based methods. The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
@@ 1 Introduction
Without the critic network, REINFORCE-based methods often struggle to accurately estimate individual tokens’ advantages. Various REINFORCE-baseline approaches have been proposed to address this limitation, but each has significant drawbacks. ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation. RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO utilizes a normalized reward across all responses. These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking. Furthermore, these methods calculate reward baselines separately for each prompt, leading to overfitting and instability on specific training prompts during optimization. Therefore, these methods require carefully curating prompt sets particular to each task.
@@ 3 Method
REINFORCE++ optimizes the PPO objective and employs the clipping strategy defined in Equation 1. To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward. Accordingly, the advantage of REINFORCE++ is the normalized reward-to-go in reinforcement learning:
@@ 4.2.2 RL from Supervised Fine-tuned Model
In real-world tasks, user prompts exhibit enormous form and intent, making it difficult to systematically assess a given model’s strengths and weaknesses. Using synthetic datasets, we can systematically manipulate key factors such as length and difficulty, allowing for a more direct and interpretable evaluation and analysis of model performance.
@@ 5 Conclusion
This paper presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences. Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global batch mean reward as a baseline to prevent overfitting and instability. Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities. Future work includes exploring adaptive normalization techniques, advanced variance reduction methods, and extending REINFORCE++ beyond RLHF settings.
@@ Appendix A Algorithm Details
Our experiments above showed that training models using multiple generated responses per prompt can further improve performance. Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation. Specifically, we sample multiple responses for each prompt and compute their average reward as the baseline to reshape the rewards. The advantage for each prompt-response pair is then defined as:
$Adv_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$
$Adv_{q,o_{t}}^{norm}=\frac{Adv_{q,o_{t}}-\text{mean}_{batch}(Adv_{q,o_{t}})}{\text{std}_{batch}(Adv_{q,o_{t}})}$
@@
REINFORCE++ shares certain similarities with PPO in its formulation. Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and employs global batch normalization as the baseline. Mathematically, this relationship is expressed as follows:
@@
where $V(s_{t})$ denotes the estimated value function at state $s_{t}$. By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term. At the same time, the introduction of global batch normalization further stabilizes training by reducing variance and preventing overfitting to individual prompts.
@@ Appendix B Acknowledgements
  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the early draft of the paper.
@@
  • Wei Shen: Led the paper writing and designed and supervised the main experiments.